# Gaudin models and moduli space of flower curves

### v1

@@ Abstract.
We introduce and study the family of trigonometric Gaudin subalgebras in $U\mathfrak{g}^{\otimes n}$ for arbitrary simple Lie algebra $\mathfrak{g}$. This is the family of commutative subalgebras of maximal possible transcendence degree that serve as a universal source for higher integrals of the trigonometric Gaudin quantum spin chain attached to $\mathfrak{g}$. We study the parameter space that indexes all possible degenerations of subalgebras from this family. In particular, we show that (rational) inhomogeneous Gaudin subalgebras of $U\mathfrak{g}^{\otimes n}$ previously studied in [FFRy] arise as certain limits of trigonometric Gaudin subalgebras. Moreover, we show that both families of commutative subalgebras glue together into the one parameterized by the space ¯$\overline{\mathcal{F}}_{n}$, which is the total space of degeneration of the Deligne-Mumford space of stable rational curves $\overline{M}_{n+2}$ to the moduli space of cactus flower curves $\overline{F}_{n}$ recently introduced in [IKLPR]. As an application, we show that trigonometric Gaudin subalgebras act on tensor products of irreducible finite-dimensional $\mathfrak{g}$-modules without multiplicities, under some explicit assumptions on the parameters in terms of two different real forms of $\overline{M}_{n+2}$. This gives rise to a monodromy action of the affine cactus group on the set of eigenstates for the trigonometric Gaudin model. We also explain the relation between the trigonometric Gaudin model and the quantum cohomology of affine Grassmannians slices.
@@ 1.1. Gaudin models
In order to explain these subalgebras, it will be necessary to introduce some notation. Fix a triangular decomposition $\mathfrak{g}=\mathfrak{n}_{+}\oplus\mathfrak{h}\oplus\mathfrak{n}_{-}$, where $\mathfrak{h}$ is a Cartan subalgebra. Fix a non-degenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $\mathfrak{g}$ and let $\Omega\in\mathfrak{g}^{\otimes 2}$ be the corresponding Casimir element. Write $\Omega=\Omega_{+}+\Omega_{0}+\Omega_{-}$ where $\Omega_{+}\in\mathfrak{n}_{+}\otimes\mathfrak{n}_{-}$, $\Omega_{0}\in\mathfrak{h}\otimes\mathfrak{h}$ and $\Omega_{-}\in\mathfrak{n}_{-}\otimes\mathfrak{n}_{+}$.
@@ 1.2.2. Verma tensor product multiplicity
image of $\text{image of }\mathcal{A}(0,z_{1}\dots,z_{n})=\text{ image of }\mathcal{A}_{\theta}^{trig}(z_{1}\dots,z_{n})$
@@ 1.3. Parameter spaces of Gaudin subalgebras
  • For the homogeneous Gaudin model, we have the parameter space $M_{n+1}:=\mathbb{C}^{n}\setminus\Delta/B$, which we identify with the moduli space of rational curves with $n+1$ distinct marked points $z_{1},\ldots,z_{n},\infty$.
  • For the trigonometric Gaudin model, we have the parameter space $M_{n+2}:=(\mathbb{C}^{\times})^{n}\setminus\Delta/\mathbb{C}^{\times}$, which we identify the moduli space of rational curves with $n+2$ distinct marked points $0,z_{1},\ldots,z_{n},\infty$.
  • For the inhomogeneous Gaudin model, we have the parameter space $F_{n}:=\mathbb{C}^{n}\setminus\Delta/\mathbb{C}$, which we identify with the moduli space of rational curves with $n$ distinct marked points $z_{1},\ldots,z_{n}$ and one distinguished point $\infty$ which caries a non-zero tangent vector.
The main goal of the present paper is to describe all possible limits of these Gaudin subalgebras and to determine the compactified parameter space for these algebras. These limit subalgebras arise as some of the parameters $z_{i}$ approach each other or tend to $\infty$ (or 00 in the trigonometric case). More precisely, we have an injective map
@@ 1.4. Relation between the models
According to Mukhin, Tarasov and Varchenko [MTV], in type A, the intermediate degenerations are related by a duality which can be seen as Howe duality on the algebras of observables and as bispectral duality on the solutions of the Bethe ansatz. From this perspective, the compactified parameter space for trigonometric Gaudin algebras from Theorem 1.2 can be compared with the compactification result for the XXX spin chain in type A, see [IR2]. For a concrete result about the degeneration of the XXX model to inhomogeneous Gaudin subalgebras, see [KMR, Theorem 8.12].
In the present paper, we prove that the trigonometric Gaudin algebras can degenerate to the inhomogeneous ones. In particular, we have the following result (Theorem 7.1):
@@
a group scheme over ${\mathbb{A}}^{1}$ which degenerates $\mathbb{C}^{\times}$ to $\mathbb{C}$. Let ${\mathcal{F}}_{n}={\mathbb{G}}^{n}\setminus\Delta/{\mathbb{G}}$. This is the total space of a degeneration of $M_{n+2}$ to $F_{n}$. We prove that for fixed $\chi\in\mathfrak{h}$, there is injective map
@@
We prove that the fibrewise (with respect to $\varepsilon\in{\mathbb{A}}^{1}$) compactification of this family is ¯$\overline{\mathcal{F}}_{n}$, another space we introduced in [IKLPR]. This is the total space of a degeneration of $\overline{M}_{n+2}$ to $\overline{F}_{n}$. Geometrically, ¯$\overline{\mathcal{F}}_{n}$ parametrizes families of marked curves where two marked points come together to form a distinguished point with a tangent vector.
@@
In the present paper, we work towards this conjecture for the trigonometric Gaudin model. We show that under certain reality conditions on the parameters, the spectrum of the trigonometric Gaudin model is simple. In particular we consider two real forms of $\overline{M}_{n+2}$: split $\overline{M}_{n+2}^{split}$ and compact $\overline{M}_{n+2}^{comp}$. (Both spaces are in fact compact; the names comes from the fact that they are related to the split (resp. compact) real forms of the torus (×$(\mathbb{C}^{\times})^{n}/\mathbb{C}^{\times}$.) On the other hand, we consider only one real form of $\overline{F}_{n}$, which we denote by $\overline{F}_{n}(\mathbb{R})$.
@@
    Let $\theta\in\mathfrak{h}^{split}$ be generic and sufficiently dominant. For every $C\in\overline{M}_{n+2}^{split}$, $\mathcal{A}_{\theta}^{trig}(C)$ acts on $V({\underline{\lambda}})$ with simple spectrum.
@@ 1.6. Monodromy of eigenvectors
In [IKLPR, Theorems 9.23, 9.24], we also proved that ¯$\overline{\mathcal{F}}^{split}_{n}$ and ¯$\overline{\mathcal{F}}^{comp}_{n}$ both deformation retract onto $\overline{F}_{n}(\mathbb{R})$ and thus the monodromy action of the fundamental groups $\pi_{1}^{S_{n}}(\overline{M}_{n+2}^{split})$, $\pi_{1}^{S_{n}}(\overline{M}_{n+2}^{comp})$ must factor through homomorphisms from these groups to $vC_{n}$.
@@
Similarly, the virtual cactus group $vC_{n}=\pi_{1}^{S_{n}}(\overline{F}_{n})$ arises in concrete coboundary monoidal categories. Namely, suppose we are given a coboundary monoidal category $\mathcal{C}$ and a faithful monoidal functor $F:\mathcal{C}\to Sets$. Then for any $n$-tuple of objects $X_{1},\ldots,X_{n}\in\mathcal{C}$, the set $F(X_{1}\otimes\ldots\otimes X_{n})=F(X_{1})\times\ldots\times F(X_{n})$ is naturally acted on by $vC_{n}$, where the symmetric group $S_{n}=\pi_{1}^{S_{n}}(pt)$ acts by naive permutations of factors (i.e. by commuters in the category of sets) and $C_{n}$ acts by commutors in the category $\mathcal{C}$.
@@
In a future work, we will use these ideas to prove the following conjectures.
@@ 1.8. Relation to quantum cohomology of slices in the affine Grassmannian
Consider the Rees algebra $U_{\hbar}\mathfrak{g}$ and the algebra $S_{\hbar}(\mathfrak{h}):=S(\mathfrak{h})\otimes\mathbb{C}[\hbar]$. Let ${\mathbf{M}}:=U_{\hbar}\mathfrak{g}\otimes_{U_{\hbar}\mathfrak{n}_{+}}\mathbb{C}[\hbar]$ be the universal Verma module, which is a $U_{\hbar}\mathfrak{g},S_{\hbar}(\mathfrak{h})$ bimodule. We will also need to consider ${\mathbf{M}}[\mu]$ which means that we shift the $S_{\hbar}(\mathfrak{h})$ action by $\mu$. We will be interested in tensor products
@@
The Gaudin algebra $\mathcal{A}(0,z_{1},\dots,z_{n})\subset U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$ acts on ${\mathbf{M}}\otimes V({\underline{\lambda}})$. This action is closely related to the trigonometric Gaudin model. More precisely, if we specialize at a generic point $\theta\in\mathfrak{h}\cong\mathfrak{h}^{*}$ and at $\hbar=1$, then there is an isomorphism (as in Proposition 5.10)
@@
The left hand side of (1) admits an action of the quantum cohomology ring $QH^{*}_{T^{\vee}\times\mathbb{C}^{\times}}(\mathrm{Gr}^{\underline{\lambda}}_{\mu})$ which depends on a quantum parameter in $H^{2}(\mathrm{Gr}^{\underline{\lambda}}_{\mu},\mathbb{C}^{\times})_{reg}$. There is a map (×$(\mathbb{C}^{\times})^{n}\setminus\Delta/\mathbb{C}^{\times}\rightarrow H^{2}(\mathrm{Gr}^{\underline{\lambda}}_{\mu},\mathbb{C}^{\times})_{reg}$ which is usually an isomorphism, so we interpret the quantum parameter as a point ${\underline{z}}\in(\mathbb{C}^{\times})^{n}\setminus\Delta/\mathbb{C}^{\times}=M_{n+2}$.
@@ 1.9. Cohomology of resolved affine Grassmannian slices
By Proposition 4.18, the corresponding limit Gaudin subalgebra $\mathcal{A}(C)$ is generated by certain diagonally embedded two point Gaudin subalgebras. However, when this algebra acts on ${\mathbf{M}}\otimes V({\underline{\lambda}})$, since all $V(\lambda_{i})$ are minuscule, the action of each of these two point Gaudin subalgebra is generated by the centre $Z=Z(U_{\hbar}\mathfrak{g})$. This leads to the following conjecture.
@@ 1.12. Organization
We thank Alexander Braverman, Pavel Etingof, Iva Halacheva, Tamas Hausel, Leo Jiang, Yu Li, Vasily Krylov, and Alexei Oblomkov for many comments and suggestions. The work of L.R. was supported by the Fondation Courtois. The work was accomplished during L.R.’s stay at Harvard University and MIT. L.R. Harvard University (especially Dennis Gaitsgory) and MIT (especially Roman Bezrukavnikov and Pavel Etingof) for their hospitality.
@@ 2.1. Simple Lie algebras
Let $\Phi_{1},\dots,\Phi_{r}$ be free generators of $S(\mathfrak{g})^{\mathfrak{g}}$, where $r=\operatorname{rk}\mathfrak{g}$. We assume that all $\Phi_{l}$ are homogeneous and set $d_{l}=\deg\Phi_{l}$. Fix a $\mathfrak{g}$-invariant non-degenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $\mathfrak{g}$ and let $\{x_{a}\},\{x^{a}\},\ a=1,\ldots,\dim\mathfrak{g}$ be dual bases of $\mathfrak{g}$ with respect to this bilinear form. Moreover, we assume that $d_{1}=2$, and that $\Phi_{1}$ is the quadratic Casimir corresponding to this bilinear form.
@@ 2.2. Set partitions and embeddings of tensor products
Throughout this paper, $n$ will be a natural number and we write $[n]:=\{1,\dots,n\}$.
For each finite set $S$, let $p(S)$ denote the set of pairs $(i,j)$ of distinct elements of $S$. We will abuse notation by abbreviating $(i,j)$ to $ij$. Similarly, we write $t(S)$ for the set of triples $(i,j,k)$ (abbreviated to $ijk$) of distinct elements of $S$.
@@
Note that when =$\mathcal{B}=\{\{1,2\}\}$, then $\Delta^{\mathcal{B}}:U\mathfrak{g}\rightarrow U\mathfrak{g}^{\otimes 2}$ is the trivial partition of $[2]$, then this is the usual coproduct. More generally, $\Delta^{\{[n]\}}:U\mathfrak{g}\rightarrow U\mathfrak{g}^{\otimes n}$ is the usual iterated coproduct, which we write simply as $\Delta$.

### v2

@@ Abstract.
We introduce and study the family of trigonometric Gaudin subalgebras in $U\mathfrak{g}^{\otimes n}$ for arbitrary simple Lie algebra $\mathfrak{g}$. This is the family of commutative subalgebras of maximal possible transcendence degree that serve as a universal source for higher integrals of the trigonometric Gaudin quantum spin chain attached to $\mathfrak{g}$. We study the parameter space that indexes all possible degenerations of subalgebras from this family. In particular, we show that (rational) inhomogeneous Gaudin subalgebras of $U\mathfrak{g}^{\otimes n}$ previously studied in [FFRy] arise as certain limits of trigonometric Gaudin subalgebras. Moreover, we show that both families of commutative subalgebras glue together into the one parameterized by the space $\overline{\mathcal{F}}_{n}$, which is the total space of degeneration of the Deligne-Mumford space of stable rational curves $\overline{M}_{n+2}$ to the moduli space of cactus flower curves $\overline{F}_{n}$ recently introduced in [IKLPR]. As an application, we show that trigonometric Gaudin subalgebras act on tensor products of irreducible finite-dimensional $\mathfrak{g}$-modules without multiplicities, under some explicit assumptions on the parameters in terms of two different real forms of $\overline{M}_{n+2}$. This gives rise to a monodromy action of the affine cactus group on the set of eigenstates for the trigonometric Gaudin model. We also explain the relation between the trigonometric Gaudin model and the quantum cohomology of minuscule resolutions of affine Grassmannian slices.
@@ 1.1. Gaudin models
In order to explain these subalgebras, it will be necessary to introduce some notation. Fix a triangular decomposition $\mathfrak{g}=\mathfrak{n}_{+}\oplus\mathfrak{h}\oplus\mathfrak{n}_{-}$, where $\mathfrak{h}$ is a Cartan subalgebra. Fix a non-degenerate invariant symmetric bilinear form $(\cdot,\cdot)$ on $\mathfrak{g}$ and let $\Omega\in\mathfrak{g}^{\otimes 2}$ be the corresponding Casimir element. Write $\Omega=\Omega_{+}+\Omega_{0}+\Omega_{-}$ where $\Omega_{+}\in\mathfrak{n}_{+}\otimes\mathfrak{n}_{-}$, $\Omega_{0}\in\mathfrak{h}\otimes\mathfrak{h}$ and $\Omega_{-}\in\mathfrak{n}_{-}\otimes\mathfrak{n}_{+}$.
@@ 1.2.2. Verma tensor product multiplicity
$\text{image of }\mathcal{A}(0,z_{1}\dots,z_{n})=\text{ image of }\mathcal{A}_{\theta}^{trig}(z_{1}\dots,z_{n})$
@@ 1.3. Parameter spaces of Gaudin subalgebras
  • For the homogeneous Gaudin model, we have the parameter space $M_{n+1}:=(\mathbb{C}^{n}\setminus\Delta)/B$, which we identify with the moduli space of rational curves with $n+1$ distinct marked points $z_{1},\ldots,z_{n},\infty$.
  • For the trigonometric Gaudin model, we have the parameter space $M_{n+2}:=((\mathbb{C}^{\times})^{n}\setminus\Delta)/\mathbb{C}^{\times}$, which we identify the moduli space of rational curves with $n+2$ distinct marked points $0,z_{1},\ldots,z_{n},\infty$.
  • For the inhomogeneous Gaudin model, we have the parameter space $F_{n}:=(\mathbb{C}^{n}\setminus\Delta)/\mathbb{C}$, which we identify with the moduli space of rational curves with $n$ distinct marked points $z_{1},\ldots,z_{n}$ and one distinguished point $\infty$ which caries a non-zero tangent vector.
The main goal of the present paper is to describe all possible limits of these Gaudin subalgebras and to determine the compactified parameter space for these algebras. These limit subalgebras arise as some of the parameters $z_{i}$ approach each other or tend to $\infty$ (or 0 in the trigonometric case). More precisely, we have an injective map
@@ 1.4. Relation between the models
[…]
In the present paper, we prove that the trigonometric Gaudin algebras can degenerate to the inhomogeneous ones. In particular, we have the following result for $\chi\in\mathfrak{h}^{reg}$ (Theorem 7.1):
@@
a group scheme over ${\mathbb{A}}^{1}$ (via the map $(z,\varepsilon)\mapsto\varepsilon$) which degenerates $\mathbb{C}^{\times}$ to $\mathbb{C}$. Let ${\mathcal{F}}_{n}=({\mathbb{G}}^{n}\setminus\Delta)/{\mathbb{G}}$. This is the total space of a degeneration of $M_{n+2}$ to $F_{n}$. We prove that for fixed $\chi\in\mathfrak{h}$, there is injective map
@@
We prove that the fibrewise (with respect to $\varepsilon\in{\mathbb{A}}^{1}$) compactification of this family is $\overline{{\mathcal{F}}}_{n}$, another space we introduced in [IKLPR]. This is the total space of a degeneration of $\overline{M}_{n+2}$ to $\overline{F}_{n}$. Geometrically, $\overline{\mathcal{F}}_{n}$ parametrizes families of marked curves where two marked points come together to form a distinguished point with a tangent vector.
@@
In the present paper, we work towards this conjecture for the trigonometric Gaudin model. We show that under certain reality conditions on the parameters, the spectrum of the trigonometric Gaudin model is simple. In particular we consider two real forms of $\overline{M}_{n+2}$: split $\overline{M}_{n+2}^{split}$ and compact $\overline{M}_{n+2}^{comp}$. (Both spaces are in fact compact; the names come from the fact that they are related to the split (resp. compact) real forms of the torus $(\mathbb{C}^{\times})^{n}/\mathbb{C}^{\times}$.) On the other hand, we consider only one real form of $\overline{F}_{n}$, which we denote by $\overline{F}_{n}(\mathbb{R})$; it appears as the degeneration of both $\overline{M}_{n+2}^{split}$ and $\overline{M}_{n+2}^{comp}$. More precisely we consider two real forms of $\overline{{\mathcal{F}}}_{n}$, $\overline{\mathcal{F}}^{split}_{n}$ and $\overline{\mathcal{F}}^{comp}_{n}$, which have general fibres $\overline{M}_{n+2}^{split}$ and $\overline{M}_{n+2}^{comp}$, respectively, and each have $\overline{F}_{n}(\mathbb{R})$ as central fibre.
@@
    Let $\theta\in\mathfrak{h}^{split}$ be generic and sufficiently dominant (i.e. $(\alpha,\theta)\gg 0$ for all positive roots $\alpha$). For every $C\in\overline{M}_{n+2}^{split}$, $\mathcal{A}_{\theta}^{trig}(C)$ acts on $V({\underline{\lambda}})$ with simple spectrum.
@@ 1.6. Monodromy of eigenvectors
In [IKLPR, Theorems 9.23, 9.24], we also proved that $\overline{\mathcal{F}}^{split}_{n}$ and $\overline{\mathcal{F}}^{comp}_{n}$ both deformation retract onto $\overline{F}_{n}(\mathbb{R})$ and thus the monodromy action of the fundamental groups $\pi_{1}^{S_{n}}(\overline{M}_{n+2}^{split})$, $\pi_{1}^{S_{n}}(\overline{M}_{n+2}^{comp})$ must factor through homomorphisms from these groups to $vC_{n}$.
@@
Similarly, the virtual cactus group $vC_{n}=\pi_{1}^{S_{n}}(\overline{F}_{n})$ arises in concrete coboundary monoidal categories [KR]. Namely, suppose we are given a coboundary monoidal category $\mathcal{C}$ and a faithful monoidal functor $F:\mathcal{C}\to Sets$. Then for any $n$-tuple of objects $X_{1},\ldots,X_{n}\in\mathcal{C}$, the set $F(X_{1}\otimes\ldots\otimes X_{n})=F(X_{1})\times\ldots\times F(X_{n})$ is naturally acted on by $vC_{n}$, where the symmetric group $S_{n}=\pi_{1}^{S_{n}}(pt)$ acts by naive permutations of factors (i.e. by commuters in the category of sets) and $C_{n}$ acts by commutors in the category $\mathcal{C}$.
@@
In a future work [KR], we will use these ideas to prove the following conjectures.
@@ 1.8. Relation to quantum cohomology of slices in the affine Grassmannian
Consider the Rees algebra $U_{\hbar}\mathfrak{g}$ and the algebra $S_{\hbar}(\mathfrak{h}):=S(\mathfrak{h})\otimes\mathbb{C}[\hbar]$. Let ${\mathbf{M}}[\mu]:=U_{\hbar}\mathfrak{g}\otimes_{U_{\hbar}\mathfrak{b}_{+}}S_{\hbar}\mathfrak{h}$ be the shifted universal Verma module, where $\mu]$ indicates that $x\in\mathfrak{h}\subset\mathfrak{b}_{+}$ acts by $x+\hbar\mu(x)$ on $S_{\hbar}(\mathfrak{h})$. This is a $U_{\hbar}\mathfrak{g},S_{\hbar}(\mathfrak{h})$-bimodule. We will be interested in tensor products
@@
The Gaudin algebra $\mathcal{A}(0,z_{1},\dots,z_{n})\subset U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$ acts on ${\mathbf{M}}\otimes V({\underline{\lambda}})$. This action is closely related to the trigonometric Gaudin model. More precisely, if we specialize at a generic point $\theta\in\mathfrak{h}\cong\mathfrak{h}^{*}$ and at $\hbar=1$, then there is an isomorphism (as in Proposition 4)
@@
The left hand side of (1) admits an action of the quantum cohomology ring $QH^{*}_{T^{\vee}\times\mathbb{C}^{\times}}(\mathrm{Gr}^{\underline{\lambda}}_{\mu})$ which depends on a quantum parameter in $H^{2}(\mathrm{Gr}^{\underline{\lambda}}_{\mu},\mathbb{C}^{\times})_{reg}$, which is the complement to the divisors defined by the Kähler roots, see [O, Section 3.1.4]. For resolved affine Grassmannian slices, this space of quantum parameters is very easy to understand, since there is a map $((\mathbb{C}^{\times})^{n}\setminus\Delta)/\mathbb{C}^{\times}\rightarrow H^{2}(\mathrm{Gr}^{\underline{\lambda}}_{\mu},\mathbb{C}^{\times})_{reg}$ which is usually an isomorphism (it is always surjective by [Da, Prop 3.4.7]), so we interpret the quantum parameter as a point ${\underline{z}}\in((\mathbb{C}^{\times})^{n}\setminus\Delta)/\mathbb{C}^{\times}=M_{n+2}$.
@@ 1.9. Cohomology of resolved affine Grassmannian slices
By Proposition 4.18, the corresponding limit Gaudin subalgebra $\mathcal{A}(C)$ is generated by certain diagonally embedded two point Gaudin subalgebras. However, when this algebra acts on ${\mathbf{M}}\otimes V({\underline{\lambda}})$, since all $V(\lambda_{i})$ are minuscule (and so, all the weight multiplicities in $V(\lambda_{i})$ are at most $1$), the tensor product of $V(\lambda_{i})$ with a generic Verma module is a multiplicity-free sum of Verma modules that are the joint eigenspaces for the diagonally embedded centre of the universal enveloping algebra. This means that the action of each of the above two point Gaudin subalgebras is generated by the diagonally embedded centre $Z=Z(U_{\hbar}\mathfrak{g})$. This leads to the following conjecture.
@@ 1.12. Organization
We thank Alexander Braverman, Pavel Etingof, Iva Halacheva, Tamas Hausel, Leo Jiang, Yu Li, Vasily Krylov, and Alexei Oblomkov for many comments and suggestions. We also thank the referee for their thoughtful feedback and constructive remarks, which helped improve the manuscript. The work of L.R. was supported by the Fondation Courtois. The work was accomplished during L.R.’s stay at Harvard University and MIT. L.R. is grateful to Harvard University (especially Dennis Gaitsgory) and MIT (especially Roman Bezrukavnikov and Pavel Etingof) for their hospitality. The research of A.I. is supported by the MSHE project No. FSMG-2024-0048.
@@ 2.1. Simple Lie algebras
Let $\Phi_{1},\dots,\Phi_{r}$ be free generators of $S(\mathfrak{g})^{\mathfrak{g}}$, where $r=\operatorname{rk}\mathfrak{g}$. We assume that all $\Phi_{l}$ are homogeneous and set $d_{l}=\deg\Phi_{l}$. Fix a $\mathfrak{g}$-invariant non-degenerate symmetric bilinear form $(\cdot,\cdot)$ on $\mathfrak{g}$ and let $\{x_{a}\},\{x^{a}\},\ a=1,\ldots,\dim\mathfrak{g}$ be dual bases of $\mathfrak{g}$ with respect to this bilinear form. Moreover, we assume that $d_{1}=2$, and that $\Phi_{1}$ is the quadratic Casimir corresponding to this bilinear form.
@@ 2.2. Set partitions and embeddings of tensor products
Throughout this paper, $n$ will be a natural number and we write $n]:=\{1,\dots,n\}$.
For each finite set $S$, let $p(S)$ denote the set of pairs $(i,j)$ of distinct elements of $S$. We will abuse notation by abbreviating $(i,j)$ to $ij$. Similarly, we write $t(S)$ for the set of triples $(i,j,k)$ (abbreviated to ijkijk) of distinct elements of $S$.
@@
Note that when $\mathcal{B}=\{\{1,2\}\}$, then $\Delta^{\mathcal{B}}:U\mathfrak{g}\rightarrow U\mathfrak{g}^{\otimes 2}$ is the trivial partition of $[2]$, then this is the usual coproduct. More generally, $\Delta^{\{[n]\}}:U\mathfrak{g}\rightarrow U\mathfrak{g}^{\otimes n}$ is the usual iterated coproduct, which we write simply as $\Delta$.